# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v3

@@ Appendix C Comparision of Methods
We provide a comprehensive comparison of different methods in Table 3. Vahdat et al. [39], Wehenkel and Louppe [43], Rombach et al. [30] perform diffusion in the latent space of a VAE to improve the efficiency of image generation. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors. Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
Recent work [25, 47, 42] has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. Cohen et al. [5] crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
@@ Appendix E Connections to Diffusion Samplers
Diffusion sampling [3, 46, 41, 45, 29, 32, 1] mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) [46], denoising diffusion sampler (DDS) [41], and time-reversed diffusion sampler (DIS) [3], which have been unified by Richter et al. [29]. Some other works [45, 1] use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. Sendera et al. [32] benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps. Overall, there are indeed some strong connections between these works and ours:
@@
  • The diffusion-structured samplers only focus on density estimation/sampling but ignore the problem of learning a generative model, which is one of the main focuses of our work. We aim to perform more accurate variational inference using an auxiliary variable model augmented by diffusion models to improve generative modeling. In our setting, $p_{\bm{\theta}}(z|x)$ is a moving target density, as we jointly learn ${\bm{\theta}}$ with ${\bm{\phi}}$, as opposed to a static target density that diffusion-structured samplers are designed to solve.
  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy [46] or off-policy [21] trajectory-based optimization. Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
  • In Equation 9, we are minimizing the forward KL divergence $\mathcal{D}_{KL}(p_{\bm{\theta}}||q_{\bm{\phi}})$, where diffusion samplers are minimizing the reverse $\mathcal{D}_{KL}(q_{\bm{\phi}}||p_{\bm{\theta}})$.
@@
[3, 46, 41, 1] [45, 29, 32, 21]
@@
Objective $D_{KL}(q||p)$ with regularizer Trajectory balance objective $D_{KL}(p||q)$ with ELBO
@@ Appendix F Priors
Pinwheel. This distribution was used in [13]. We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in [22]. For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
@@ Appendix G Model Architecture
All methods use the same architecture for encoder $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ and decoder $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$, excluding the extra parts specific to each method which we describe below, for the same dataset. For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units. For CIFAR-10, the encoder is a 4-layer convolutional neural network with (16, 32, 64, 128) channels with a linear layer on top, and the decoder is a 4-layer tranposed convolutional neural network with (64, 32, 16, 3) channels where a linear layer is used to first turn the feature dimension from 2 to 64.
@@ Appendix H Training Details
For training, we update the parameters for each batch of inputs by alternating between the ELBO phase (optimizing ${\bm{\theta}}$ and ${\bm{\phi}}$ with respect to the ELBO, i.e., the reconstruction term and the prior matching term) and the sleep phase (optimizing ${\bm{\phi}}$ with respect to the sleep term). We use Adam optimizer and latent size of 2 for all of our experiments. Each algorithm takes roughly 2 hours on a single Nvidia GeForce RTX 3090 to complete one run of experiment. The training details of each algorithm are detailed below:
@@ Appendix I Genotype Analysis Experiments Details
Method Pinwheel Swiss Roll Square ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$ AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\bm{-2.10\pm 0.31}$ $1.71\pm 0.77$ AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\bm{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$ H-AEVB $-4.42\pm 0.46$ $\bm{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$ PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$ DDVI $\bm{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\bm{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\bm{1.58\pm 0.09}$
@@
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\bm{-1.60\pm 0.29}$ $\bm{0.49\pm 0.01}$ $\bm{1.09\pm 0.05}$ $\bm{-4.13\pm 1.51}$ $\bm{0.47\pm 0.09}$ $\bm{2.29\pm 0.08}$ $\bm{-1.73\pm 0.64}$ $\bm{0.49\pm 0.01}$ $\bm{1.48\pm 0.02}$
@@
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\bm{1.27\pm 0.21}$ $\bm{3.86\pm 1.17}$ $\bm{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
@@
Appendix J Data Licenses
Yann LeCun and Corinna Cortes hold the copyright of MNIST dataset, which is a derivative work from original NIST datasets. CIFAR-10 is under MIT license. 1000 Genomes dataset is under the creative commons Attribution-NonCommercial-ShareAlike 3.0 Unported licence.

### v4

@@ Appendix C Comparision of Methods
We provide a comprehensive comparison of different methods in Table 4. Vahdat, Kreis, and Kautz (2021); Wehenkel and Louppe (2021); Rombach et al. (2022) perform diffusion in the latent space of a VAE to improve the efficiency of image generation. Their goal is high sample quality, and they introduce into $p$ hierarchical latents with simple Gaussian priors. Our goal is different: we seek a method to fit a $p$ with structured latents (e.g., in probabilistic programming or in science applications, users introduce prior knowledge via hand-crafted $p$), and we improve variational inference in this structured model by introducing auxiliary latents into $q$.
Recent work (Preechakul et al. 2022; Zhang, Zhao, and Lin 2022; Wang et al. 2023) has also melded auto-encoders with diffusion models, focusing on semantically meaningful low-dimensional latents in a diffuser $p$. Cohen et al. (2022) crafts a diffusion bridge linking a continuous coded vector to a non-informative prior distribution.
@@ Appendix E Connections to Diffusion Samplers
Diffusion sampling (Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Zhang et al. 2023; Richter, Berner, and Liu 2023; Sendera et al. 2024; Akhound-Sadegh et al. 2024) mainly focuses on the task of drawing samples from unnormalized distributions and estimating the partition function. These works draw connections between diffusion (learning the denoising process) and stochastic control (learning the Föllmer drift), leading to several approaches, e.g., path integral sampler (PIS) (Zhang and Chen 2021), denoising diffusion sampler (DDS) (Vargas, Grathwohl, and Doucet 2023), and time-reversed diffusion sampler (DIS) (Berner, Richter, and Ullrich 2022), which have been unified by Richter, Berner, and Liu (2023). Some other works (Zhang et al. 2023; Akhound-Sadegh et al. 2024) use continuous generative flow networks (GFlowNets) – deep reinforcement learning algorithms adapted to variational inference that offers stable off-policy training and thus flexible exploration. Sendera et al. (2024) benchmarked these previous diffusion-structured amortized inference methods and studied how to improve credit assignment in diffusion samplers, which refers to the propagation of learning signals from the target density to the parameters of earlier sampling steps. Overall, there are indeed some strong connections between these works and ours:
@@
  • The diffusion-structured samplers only focus on density estimation/sampling but ignore the problem of learning a generative model, which is one of the main focuses of our work. We aim to perform more accurate variational inference using an auxiliary variable model augmented by diffusion models to improve generative modeling. In our setting, $p_{\boldsymbol{\theta}}(z|x)$ is a moving target density, as we jointly learn ${\boldsymbol{\theta}}$ with ${\boldsymbol{\phi}}$, as opposed to a static target density that diffusion-structured samplers are designed to solve.
  • To tackle the challenge of credit assignment – propagating weak learning signals through the sampling trajectory, the techniques proposed in diffusion-structured samplers are mostly based on partial trajectory information, which has higher training costs over on-policy (Zhang and Chen 2021) or off-policy (Malkin et al. 2022) trajectory-based optimization. Instead, we introduce a wake-sleep optimization algorithm and its simplified version to alleviate the weak learning signal issue and optimize the evidence lower bound in a better way.
  • In Equation 8, we are minimizing the forward KL divergence $\text{KL}(p_{\boldsymbol{\theta}}||q_{\boldsymbol{\phi}})$, where diffusion samplers are minimizing the reverse $\text{KL}(q_{\boldsymbol{\phi}}||p_{\boldsymbol{\theta}})$.
@@
(Berner, Richter, and Ullrich 2022; Zhang and Chen 2021; Vargas, Grathwohl, and Doucet 2023; Akhound-Sadegh et al. 2024) (Zhang et al. 2023; Richter, Berner, and Liu 2023; Sendera et al. 2024; Malkin et al. 2022)
@@
Objective $\text{KL}(q||p)$ with regularizer Trajectory balance objective $\text{KL}(p||q)$ with ELBO
@@ Appendix F Priors
Pinwheel. This distribution was used in (Johnson et al. 2016). We define the number of clusters to be 10. For semi-supervised learning experiments, this prior is partitioned into 10 partitions, each partition being a cluster.
Swiss Roll. This distribution was used in (Marsland 2014). For semi-supervised learning experiments, this prior is partitioned into 10 partitions. The samples from the prior can actually be characterized by a single scalar representing how far you are long the swiss roll from the center. The paritioning is done by creating 10 equal-length intervals in this 1D space.
@@ Appendix G Model Architecture
All methods use the same architecture for encoder $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})$ and decoder $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$, excluding the extra parts specific to each method which we describe below, for the same dataset. For MNIST, the encoder and decoder are multi-layer perceptron with two hidden layers, each with 1000 hidden units. For CIFAR-10, the encoder is a 4-layer convolutional neural network with (16, 32, 64, 128) channels with a linear layer on top, and the decoder is a 4-layer tranposed convolutional neural network with (64, 32, 16, 3) channels where a linear layer is used to first turn the feature dimension from 2 to 64.
@@ Appendix H Training Details
For training, we update the parameters for each batch of inputs by alternating between the ELBO phase (optimizing ${\boldsymbol{\theta}}$ and ${\boldsymbol{\phi}}$ with respect to the ELBO, i.e., the reconstruction term and the prior matching term) and the sleep phase (optimizing ${\boldsymbol{\phi}}$ with respect to the sleep term). We use Adam optimizer and latent size of 2 for all of our experiments. Each algorithm takes roughly 2 hours on a single Nvidia GeForce RTX 3090 to complete one run of experiment. The training details of each algorithm are detailed below:
@@ Appendix I Genotype Analysis Experiments Details
Method Pinwheel Swiss Roll Square
ELBO Latent NLL ELBO Latent NLL ELBO Latent NLL
AEVB $-12.96\pm 1.81$ $3.26\pm 0.60$ $-12.87\pm 4.55$ $6.25\pm 1.58$ $-7.91\pm 0.11$ $2.91\pm 0.17$
AEVB-IAF $-3.24\pm 0.16$ $1.71\pm 0.84$ $-4.03\pm 0.73$ $5.51\pm 0.51$ $\mathbf{-2.10\pm 0.31}$ $1.71\pm 0.77$
AAEB $N/A$ $1.70\pm 0.41$ $N/A$ $\mathbf{3.18\pm 0.22}$ $N/A$ $1.67\pm 0.17$
H-AEVB $-4.42\pm 0.46$ $\mathbf{1.69\pm 0.17}$ $-5.36\pm 0.77$ $5.74\pm 0.55$ $-2.86\pm 0.11$ $1.64\pm 0.09$
PIS ${-2.92\pm 1.23}$ ${3.61\pm 0.62}$ $-4.14\pm 0.49$ $7.14\pm 0.14$ $-4.85\pm 0.06$ $3.91\pm 0.06$
DDVI $\mathbf{-1.38\pm 0.44}$ ${1.75\pm 0.53}$ $\mathbf{-3.05\pm 0.65}$ ${5.66\pm 2.63}$ $-2.47\pm 0.30$ $\mathbf{1.58\pm 0.09}$
@@
Method Pinwheel Swiss Roll Square ELBO Acc Latent NLL ELBO Acc Latent NLL ELBO Acc Latent NLL AEVB $-17.14\pm 1.46$ $0.30\pm 0.05$ $2.32\pm 0.27$ $-17.89\pm 5.21$ $0.20\pm 0.07$ $6.56\pm 2.25$ $-13.30\pm 1.50$ $0.30\pm 0.05$ $1.95\pm 0.28$ AEVB-IAF $-5.70\pm 0.07$ $0.47\pm 0.01$ $1.62\pm 0.05$ $-5.53\pm 2.82$ $0.28\pm 0.08$ $6.82\pm 1.90$ $-4.41\pm 0.53$ $0.36\pm 0.01$ $1.58\pm 0.15$ AAEB $N/A$ $0.25\pm 0.01$ $1.77\pm 0.14$ $N/A$ $0.23\pm 0.01$ $3.38\pm 0.30$ $N/A$ $0.23\pm 0.04$ $1.74\pm 0.15$ DDVI $\mathbf{-1.60\pm 0.29}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.09\pm 0.05}$ $\mathbf{-4.13\pm 1.51}$ $\mathbf{0.47\pm 0.09}$ $\mathbf{2.29\pm 0.08}$ $\mathbf{-1.73\pm 0.64}$ $\mathbf{0.49\pm 0.01}$ $\mathbf{1.48\pm 0.02}$
@@
Method Latent NLL - Pinwheel Latent NLL - Swiss Roll Latent NLL - Square AEVB $1.68\pm 0.31$ $5.65\pm 1.58$ $2.78\pm 0.61$ AEVB-IAF $1.64\pm 0.73$ $4.43\pm 1.09$ $1.68\pm 0.27$ AAEB -- -- -- H-AEVB $2.25\pm 3.02$ $4.04\pm 4.62$ $2.22\pm 2.03$ DDVI $\mathbf{1.27\pm 0.21}$ $\mathbf{3.86\pm 1.17}$ $\mathbf{1.56\pm 0.09}$ DDVI (w/o sleep term) $2.12$ $5.25$ $2.97$
@@
Appendix J ELBO for Auxiliary-Variable Generative Models
We aim to derive the lower bound for the log-likelihood $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ by introducing auxiliary variables and applying the Evidence Lower Bound (ELBO) twice.
Introduce the latent variable $\mathbf{z}$ and apply the ELBO:
[…]
$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (17)
Then, introduce an auxiliary variable $\mathbf{y}$ and apply the ELBO again:
$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\right)$ (19)
$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\log p(\mathbf{z})$ (20)
$\geq\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]\right]-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{x},\mathbf{z})}\right]\right]+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}\left[\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})}\left[\log p(\mathbf{z})\right]\right]$ (21)
$=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})\right]-\text{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})\,\|\,r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z})\right)$ (22)
Appendix K Diffusion Regularization
We begin with the definition of $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, derived as a lower bound on $\mathcal{L}_{\text{sleep}}({\boldsymbol{\phi}})$:
$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\left[\mathbb{E}_{r}\left[\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}\right]\right]+\bar{H}(p_{\boldsymbol{\theta}}).$
To simplify $\mathcal{L}_{\text{diff}}$, we leverage the Markov structure of the forward process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ and the reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$.
The forward process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ is decomposed as:
$r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x}),$
where $\mathbf{y}_{0}=\mathbf{z}$.
The reverse process $q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ is decomposed as:
$q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x}),$
where $\mathbf{y}_{0}=\mathbf{z}$ and $\mathbf{y}_{1:T}$ are increasingly noisy versions of $\mathbf{z}$.
Substituting the factorizations of $r$ and $q_{\boldsymbol{\phi}}$ into the definition of $\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$, and rewriting the logarithm and rearranging terms, we have:
$\mathcal{L}_{\text{diff}}({\boldsymbol{\phi}})$ $=\mathbb{E}_{p_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{z})}\Bigg{[}\mathbb{E}_{r(\mathbf{y}_{1:T}|\mathbf{z},\mathbf{x})}\Bigg{[}\log\frac{q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})}{\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})}\Bigg{]}\Bigg{]}+\bar{H}(p_{\boldsymbol{\theta}})$
$=\mathbb{E}_{r(\mathbf{y}_{1:T},\mathbf{z},\mathbf{x})}\left[\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}\text{KL}(r_{t}||q_{t})\right]-\text{KL}(r(\mathbf{y}_{T}|\mathbf{z},\mathbf{x})||q_{\boldsymbol{\phi}}(\mathbf{y}_{T}|\mathbf{x}))+\bar{H}(p_{\boldsymbol{\theta}})$
Here, $\log q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})$ corresponds to the reconstruction term for the initial latent state $\mathbf{z}$, while the summation represents the KL divergence between the forward and reverse processes at each intermediate step. $r_{t}=r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{z},\mathbf{x})$ is the conditional forward distribution at step $t$, $q_{t}=q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ is the reverse process distribution at step $t$. $\bar{H}(p_{\boldsymbol{\theta}})$ is the expected conditional entropy of $p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\boldsymbol{\phi}}$.
This form mirrors the ELBO derivation for diffusion models (Sohl-Dickstein et al. 2015), where each step in the Markov chain contributes a KL divergence term, and the reconstruction term arises from the connection between the noisy latent variables and the original data.